How much water must be added to 30oz of a 80% solution to change it to a 60% solution?
Answer by lwsshak3(11628)

How much water must be added to 30oz of a 80% solution to change it to a 60% solution?
***
let x=amt of water to add
80%*30+0%x=60%(x+30)
24+0=0.6x+18
0.6x=24-18=6
x=10
amt of water to add=10 oz